a cylinder of mass 10kg is sliding on a plane with an initial velocity of 10m/s . if the coefficient of friction between surface and the cylinder is 0.5 , then before stopping it will travel?

Grade:12th pass

2 Answers

Kumari Sanjana
6 years ago
Using 3rd equation of motion v^2-u^2=2axV=0m/sU=10m/sa=-0.5*g=-5m/s^2We can find x-100=-2*5*xx=10m
Prashant
6 years ago
Initial speed=10m/s,Final speed=0,,,Acc[=mue*g=•5×10=5m/s^2,,,Applying third equation,,We get distance as 10 m